OpenStudy (anonymous):

If the terminal side of the angle θ is on the line y=- 1/4x and is in quadrant II find the sinθ and cotθ. #HELPPP :/

OpenStudy (e.mccormick):

Did you start with a sketch?

OpenStudy (anonymous):

no.

OpenStudy (e.mccormick):

It can help you find the reference angle you need by making a right triangle.

OpenStudy (anonymous):

thanks.

OpenStudy (ikram002p):

if y =(-1/4)x then tanθ=-1/4 , So the angle would be tan^-1(-1/4)=14, since its in quadrant II sinθ is positive =0.24 and cosθ is negative = -0.97

OpenStudy (e.mccormick):

https://www.desmos.com/calculator/f4ii9pdkrh With that triangle, or another on the same line, I can find the trig ratios for the angles. Then 180- the angle from the origin down to the x axis is the measure of the one you want.

OpenStudy (e.mccormick):

Oops ...down from the line that goes through the origin....

OpenStudy (e.mccormick):

Which is basically the same thing ikram002p said. He is just using the slope alone, which is what gives me the ratio of the sides as well.